OpenStudy (anonymous):

Scientific notation: 4.6 x 10^-7/9.2 x 10^-9 If there's work to be shown, please show b/c that would help. :)

OpenStudy (anonymous):

first divide 4.6 bu 9.2 = 0.5 and 10^-7 / 10^-9 = 10 ^2 so quotient = 0.5 * 10^2 in scientific notation this is 5 * 10^1 or 5 * 10 scientific notation must be of the form n.nnn.. * 10^n

OpenStudy (anonymous):

ok with that?

OpenStudy (anonymous):

So the answer is 5 x 10? or .5 x 10^2?

OpenStudy (anonymous):

no the answer is 5 * 10 . 0.5* 10^2 is not scientific notation